### v3

# Dp-finite and Noetherian NIP integral domains

Abstract
We prove some results on NIP integral domains, especially those that are Noetherian or have finite dp-rank. If $R$ is an NIP Noetherian domain that is not a field, then $R$ is a semilocal ring of Krull dimension 1, and the fraction field of $R$ has characteristic 0. Assuming the henselianity conjecture (on NIP valued fields), $R$ is a henselian local ring. Additionally, we show that integral domains of finite dp-rank are henselian local rings. Finally, we lay some groundwork for the study of Noetherian domains of finite dp-rank, and we classify dp-minimal Noetherian domains.
00footnotetext: 2020 Mathematical Subject Classification: 03C60, 12L12. Key words and phrases: NIP, dp-rank, Noetherian rings
1 Introduction
In this paper, rings are commutative and unital. We consider rings and fields as structures in a language $\mathcal{L}$ expanding the language of rings. “Definable” means “definable with parameters”. Recall that a local ring is a ring $R$ with a unique maximal ideal $\mathfrak{m}$, and a local ring $R$ is henselian if any polynomial $c_{n}x^{n}+\cdots+c_{1}x+c_{0}\in R[x]$ with $c_{0}\in\mathfrak{m}$ and $c_{1}\notin\mathfrak{m}$ has a root in $\mathfrak{m}$. A valuation ring is an integral domain $\mathcal{O}$ such that for every non-zero $x$ in the fraction field $\operatorname{Frac}(\mathcal{O})$, at least one of $x$ or $x^{-1}$ is in $\mathcal{O}$. Valuation rings are always local rings.
1.1 Henselianity
The following henselianity conjecture is part of the conjectural classification of NIP fields discussed in [2, 13, 14].
Conjecture 1.1 (Henselianity conjecture).
If $\mathcal{O}$ is an NIP valuation ring, then $\mathcal{O}$ is henselian.
Conjecture 1.1 is known to hold in the following cases:
  1. 1.
    $\operatorname{Frac}(\mathcal{O})$ has characteristic $p>0$, or equivalently, $\mathcal{O}$ is an $\mathbb{F}_{p}$-algebra [20, Theorem 2.8].
  2. 2.
    $\mathcal{O}$ is dp-finite [19, Corollary 4.16(3)], meaning that the dp-rank of $\mathcal{O}$ is finite.
See [31, Section 4.2] for information on dp-rank. Note that “NIP” is equivalent to “bounded dp-rank”, and “dp-finite” is a proper subclass of “NIP”. Dp-minimal means “dp-rank 1”. By [6, Proposition 2.8(2)], a valuation ring $\mathcal{O}$ has the same dp-rank as the corresponding valued field, i.e., the structure $(K,+,\cdot,\mathcal{O})$ where $K=\operatorname{Frac}(\mathcal{O})$.
We propose the following generalization of Conjecture 1.1.
Conjecture 1.2 (Generalized henselianity conjecture).
If $R$ is an NIP ring, then $R$ is a direct product of finitely many henselian local rings. In particular, if $R$ is an NIP integral domain, then $R$ is a henselian local ring.
See Proposition 3.6 for some equivalent forms of this conjecture. When $R$ is an $\mathbb{F}_{p}$-algebra, Conjecture 1.2 holds by [22, Theorem 3.21]. In this paper, we verify the dp-finite case:
Theorem 1.3 (= Theorem 5.9).
If $R$ is a dp-finite ring, then $R$ satisfies Conjecture 1.2: $R$ is a direct product of finitely many henselian local rings.
We also show that the henselianity conjecture implies Conjecture 1.2 in certain cases, namely Theorems 1.4 and 1.6 below.
Theorem 1.4 (= Theorem 7.9).
Assume the henselianity conjecture. If $R$ is a Noetherian NIP ring, then $R$ satisfies Conjecture 1.2: $R$ is a direct product of finitely many henselian local rings.
We will say more about NIP Noetherian rings in Section 1.2 below. Underlying Theorems 1.3 and 1.4 is a result on $W_{n}$-rings.”
Definition 1.5.
Fix $n\geq 1$. A ring $R$ is a $W_{n}$-ring if the following condition holds: if $S$ is a finite subset of $R$, then there is a subset $S^{\prime}\subseteq S$ with $|S^{\prime}|\leq n$ such that $S^{\prime}$ and $S$ generate the same ideal. A $W_{n}$-domain is a $W_{n}$-ring that is an integral domain.
For example, a $W_{1}$-domain is the same thing as a valuation ring. We give several equivalent characterizations of $W_{n}$-rings in Section 2. Note that our terminology is slightly different from [18, Definition 2.5], where $W_{n}$-ring” meant $W_{n}$-domain”. [^1] As an example of the difference, $\mathbb{Z}/4\mathbb{Z}$ is a $W_{1}$-ring but not a $W_{1}$-domain.
Theorem 1.6 (= Theorem 4.10).
Assume the henselianity conjecture. If $R$ is an NIP $W_{n}$-ring, then $R$ satisfies Conjecture 1.2: $R$ is a direct product of finitely many henselian local rings.
The connection between dp-finiteness, Noetherianity, and $W_{n}$-rings is given as follows:
Lemma 1.7 ($\subseteq$ Corollary 7.6 $\cup$ Lemma 5.6).
Let $R$ be an NIP ring.
  1. 1.
    If $R$ is Noetherian, then $R$ is a $W_{n}$-ring for some $n$.
  2. 2.
    If $R$ is dp-finite and $R/\mathfrak{m}$ is infinite for every maximal ideal $\mathfrak{m}\subseteq R$, then $R$ is a $W_{n}$-ring for $n=\operatorname{dp-rk}(R)$.
Remark 1.8.
Theorems 1.3, 1.4, and 1.6 suggest the following natural question: does the henselianity conjecture (Conjecture 1.1) imply Conjecture 1.2? I could not see how to prove this. A more promising question is whether the Shelah conjecture implies Conjecture 1.2. The Shelah conjecture says that if $K$ is an NIP field, then one of the following holds: $K$ is finite, $K$ is algebraically closed, $K$ is real closed, or $K$ admits a non-trivial henselian valuation. Like the henselianity conjecture, the Shelah conjecture is part of the conjectural classification of NIP fields. In fact, the Shelah conjecture implies a complete classification of NIP fields [2, Theorem 7.1]. By work of Halevi, Hasson, and Jahnke [14], the Shelah conjecture implies the henselianity conjecture. Perhaps similar arguments could be used to show that the Shelah conjecture implies Conjecture 1.2. Proposition 3.6 gives some alternate formulations of Conjecture 1.2 which may be useful for this purpose.
Remark 1.9.
Conjecture 1.2 is a statement about NIP rings, but it also has applications for definable field topologies on NIP fields (see Section 5 for the definition of “definable topology”). In a later paper [24], we will show that Conjecture 1.2 implies the following statement:
Let $(K,+,\cdot,\ldots)$ be an NIP expansion of a field, and $\tau$ be a definable field topology on $K$. Then $K$ is large in the sense of Pop [28] and $\tau$ is “generalized topologically henselian” in the sense of [9, Definition 8.1] (essentially meaning that the inverse function theorem holds for polynomial maps).
In particular, this statement holds when $\operatorname{char}(K)>0$ or $K$ is dp-finite.
1.2 NIP Noetherian domains
Another goal of this paper is to initiate the study of NIP Noetherian domains. If we want a classification of NIP Noetherian domains, we should either work modulo the conjectural classification of NIP fields (see [2, Theorem 7.1]), or focus on cases of low dp-rank such as dp-minimal and dp-finite rings.
There are several motivations for studying NIP Noetherian domains. For model theorists, the problem of classifying NIP Noetherian domains can be seen as a first step beyond classifying NIP fields, towards the harder problem of classifying NIP integral domains. Additionally, a classification of NIP Noetherian domains may provide new examples of NIP theories. For commutative algebraists, the class of Noetherian domains has central importance and one would like to know which Noetherian domains are amenable to the tools of model theory.
In this paper, we take some first steps towards the classification of NIP Noetherian domains, especially dp-finite Noetherian domains. Here are the main results. Recall that the Krull dimension of a ring $R$ is the supremum of lengths of finite chains of prime ideals $\mathfrak{p}_{0}\subsetneq\mathfrak{p}_{1}\subsetneq\cdots\subsetneq\mathfrak{p}_{n}$ in $R$.
Theorem 1.10 ($\subseteq$ Theorem 7.8 $\cup$ Corollary 7.6).
Let $R$ be an NIP Noetherian domain. Suppose $R$ is not a field.
  1. 1.
    $\operatorname{Frac}(R)$ has characteristic 0.
  2. 2.
    $R$ has finitely many maximal ideals $\mathfrak{m}_{1},\ldots,\mathfrak{m}_{n}$.
  3. 3.
    $R$ has Krull dimension 1, and so $\{0,\mathfrak{m}_{1},\ldots,\mathfrak{m}_{n}\}$ are the only prime ideals of $R$.
In the dp-finite case, we get the following trichotomy:
Theorem 1.11 (= Theorem 8.2).
Let $R$ be a dp-finite Noetherian domain. Then $R$ is a local ring and one of three things happens:
  1. 1.
    $R$ is not a field. $\operatorname{Frac}(R)$ and the residue field both have characteristic 0.
  2. 2.
    $R$ is not a field. $\operatorname{Frac}(R)$ has characteristic 0 and the residue field is finite.
  3. 3.
    $R$ is a field.
Lastly, we get a classification of dp-minimal Noetherian domains, building off the known classifications of dp-minimal fields and dp-minimal valuation rings. Recall that a structure is dp-minimal if it has dp-rank 1. See [23, Theorem 1.3] for the classification of dp-minimal fields, which is too complicated to state here. Recall that a discrete valuation ring (DVR) is a valuation ring $\mathcal{O}$ with fraction field $K$ such that the corresponding value group $K^{\times}/\mathcal{O}^{\times}$ is isomorphic to $\mathbb{Z}$. By [23, Theorems 1.5–1.6], dp-minimal DVRs are classified as follows:  [^2] Recall that a valuation ring $(\mathcal{O},+,\cdot)$ has the same dp-rank as its corresponding valued field $(\operatorname{Frac}(\mathcal{O}),+,\cdot,\mathcal{O})$ by [6, Proposition 2.8(2)], so the reference to [23, Theorems 1.5–1.6] is relevant.
Fact 1.12.
Let $R$ be a DVR. Then $R$ is dp-minimal iff $R$ is henselian, $\operatorname{Frac}(R)$ has characteristic 0, and one of the following holds:
  1. 1.
    The residue field is a dp-minimal field of characteristic 0.
  2. 2.
    The residue field is finite.
We call these two cases the “equicharacteristic” and “mixed characteristic” cases of dp-minimal DVRs. If $R$ is an equicharacteristic dp-minimal DVR, then $R\equiv K[[t]]$ for some dp-minimal field $K$ of characteristic 0, by the Ax-Kochen-Ershov principle. If $R$ is a mixed characteristic dp-minimal DVR, then $R\equiv\mathcal{O}_{K}$, where $K$ is a non-archimedean local field of characteristic 0 (i.e., a finite extension of $\mathbb{Q}_{p}$), and $\mathcal{O}_{K}$ is its ring of integers. [^3] Indeed, the valued field $\operatorname{Frac}(R)$ is henselian with finite residue field and value group $\mathbb{Z}$, so it is $p$-adically closed by [29, Theorem 3.1], hence elementarily equivalent to an extension of $\mathbb{Q}_{p}$ by [13, Fact 3.7].
Theorem 1.13 (= Theorem 8.9).
The following is a complete list of dp-minimal Noetherian domains:
  1. 1.
    Equicharacteristic dp-minimal DVRs.
  2. 2.
    Finite index subrings of mixed characteristic dp-minimal DVRs.
  3. 3.
    Dp-minimal fields.
Here, a “finite index” subring of $R$ is a subring $R_{0}\subseteq R$ with finite index in $(R,+)$.
In subsequent work with d’Elbée and Halevi [7], Theorem 1.13 is generalized to a complete classification of dp-minimal integral domains.
2 Breadth in modular lattices and modules
In this section, we review some facts about breadth of lattices, modules, and rings, from [21, 16, 17, 18].
Remark 2.1.
Breadth was called “reduced rank” in [16, 17] and earlier drafts of [21] and called “cube rank” or “weight” in [18]. Yatir Halevi pointed out that “breadth” is the standard terminology in lattice theory [4, Exercise II.5.6]. It seems reasonable to extend this terminology from lattices to modules and rings.
In the following, a “lattice” means an unbounded lattice, i.e., a partial order $(\Lambda,\leq)$ in which every finite non-empty subset has a infimum and supremum. Unbounded lattices can be regarded as algebraic structures $(\Lambda,\wedge,\vee)$ satisfying certain identities [4, Section I.5], where $\wedge$ and $\vee$ are the binary infimum and supremum, respectively. In particular, a “sublattice” means a subset $\Lambda_{0}\subseteq\Lambda$ closed under $\wedge$ and $\vee$, but not necessarily containing $\max(\Lambda)$ or $\min(\Lambda)$ when they exist.
Recall that a lattice $\Lambda$ is modular [4, Section I.7] if it satisfies the modular identity:
$(x\vee a)\wedge b=(x\wedge b)\vee a\text{ when }a\leq b.$
For any $R$-module $M$, the lattice $\operatorname{Sub}_{R}(M)$ of $R$-submodules of $M$ is modular [4, Theorem VII.1].
Let $\Lambda$ be a non-empty modular lattice, such as the lattice $\operatorname{Sub}_{R}(M)$. The length of $\Lambda$, written $\ell(\Lambda)$, is the largest $n$ such that there is a chain $x_{0}<x_{1}<\cdots<x_{n}$ in $\Lambda$, or $\infty$ if no maximum exists. The breadth of $\Lambda$, written $\operatorname{br}(\Lambda)$, is an element of $\{0,1,2,\ldots\}\cup\{\infty\}$ characterized in one of the following ways:
  1. 1.
    $\operatorname{br}(\Lambda)\geq n$ if there is a sublattice of $\Lambda$ isomorphic to the powerset of $n$.
  2. 2.
    $\operatorname{br}(\Lambda)\leq n$ if for any $x_{1},\ldots,x_{n+1}\in\Lambda$, there is some $i$ such that
    $x_{1}\vee\cdots\vee x_{n+1}=x_{1}\vee\cdots\vee\widehat{x_{i}}\vee\cdots\vee x_{n+1},$
    where the hat indicates omission.
  3. 3.
    $\operatorname{br}(\Lambda)\leq n$ if for any $x_{1},\ldots,x_{n+1}\in\Lambda$, there is some $i$ such that
    $x_{1}\wedge\cdots\wedge x_{n+1}=x_{1}\wedge\cdots\wedge\widehat{x_{i}}\wedge\cdots\wedge x_{n+1}.$
These are equivalent by [21, Lemma 9.9]. If $R$ is a ring and $M$ is an $R$-module, the length of $M$, written $\ell(M)$ or $\ell_{R}(M)$, is the length of the lattice $\operatorname{Sub}_{R}(M)$. Equivalently, $\ell(M)$ is the largest $n$ such that there exists a chain of submodules $M_{0}\subsetneq M_{1}\subsetneq\cdots\subsetneq M_{n}\subseteq M$. The breadth of $M$, written $\operatorname{br}(M)$ or $\operatorname{br}_{R}(M)$, is the breadth of the lattice $\operatorname{Sub}_{R}(M)$.
Remark 2.2.
By [16, Remark 6.7] and [17, Proposition 7.3], we can equivalently characterize breadth as follows:
  1. 1.
    $\operatorname{br}(M)\geq n$ iff there are submodules $N\leq N^{\prime}\leq M$ such that $N^{\prime}/N$ is a direct sum of $n$ non-trivial $R$-modules.
  2. 2.
    $\operatorname{br}(M)\leq n$ iff for any $x_{1},\ldots,x_{n+1}\in M$, there is $i$ such that
    $Rx_{1}+\cdots+Rx_{n+1}=Rx_{1}+\cdots+\widehat{Rx_{i}}+\cdots+Rx_{n+1}.$
Fact 2.3 ([18, Lemma 2.2]).
If $M$ is an $R$-module and $R^{\prime}$ is a subring of $R$, then $\operatorname{br}_{R}(M)\leq\operatorname{br}_{R^{\prime}}(M)$.
Remark 2.4.
Similarly, if $I$ is an ideal in $R$ and $M$ is an $R/I$-module, regarded as an $R$-module in the natural way, then $\operatorname{br}_{R}(M)=\operatorname{br}_{R/I}(M)$, because the lattices of submodules are the same:
$\operatorname{Sub}_{R}(M)=\operatorname{Sub}_{R/I}(M).$
Remark 2.5.
If $V$ is a vector space over a field $K$, then $\operatorname{br}_{K}(V)=\ell_{K}(V)=\dim_{K}(V)$. This follows easily using the definitions, or using the following facts.
Fact 2.6 ([16, Proposition 6.9]).
Let $R$ be a ring and $0\to M_{1}\to M_{2}\to M_{3}\to 0$ be a short exact sequence of $R$-modules. Then the following hold:
  1. 1.
    $\operatorname{br}(M_{1})\leq\operatorname{br}(M_{2})$ and $\operatorname{br}(M_{3})\leq\operatorname{br}(M_{2})$.
  2. 2.
    $\operatorname{br}(M_{2})\leq\operatorname{br}(M_{1})+\operatorname{br}(M_{3})$.
  3. 3.
    If the sequence splits (i.e., $M_{2}\cong M_{1}\oplus M_{3}$), then $\operatorname{br}(M_{2})=\operatorname{br}(M_{1})+\operatorname{br}(M_{3})$.
Fact 2.6 is analogous to the classic Jordan-Hölder theorem for length:
Fact 2.7 ([11, Theorem 2.13(a)]).
If $0\to M_{1}\to M_{2}\to M_{3}\to 0$ is a short exact sequence of $R$-modules, then $\ell(M_{2})=\ell(M_{1})+\ell(M_{3})$.
If $R$ is a ring, the breadth $\operatorname{br}(R)$ is the breadth of $R$ as an $R$-module, i.e., the breadth of the lattice of ideals. When $R$ is an integral domain with fraction field $K$, we have $\operatorname{br}(R)=\operatorname{br}_{R}(K)$ by [18, Lemma 2.4]. The $W_{n}$-rings of Definition 1.5 are precisely the rings with $\operatorname{br}(R)\leq n$, by Remark 2.2. Valuation rings are the same thing as $W_{1}$-domains [18, Proposition 2.6].
A multivaluation ring is a finite intersection of valuation rings. Multivaluation rings are examples of $W_{n}$-domains:
Fact 2.8.
Let $\mathcal{O}_{1},\ldots,\mathcal{O}_{n}$ be pairwise incomparable non-trivial valuation rings on a field $K$, and let $R=\bigcap_{i=1}^{n}\mathcal{O}_{i}$.
  1. 1.
    $R$ has exactly $n$ maximal ideals $\mathfrak{m}_{1},\ldots,\mathfrak{m}_{n}$, and $\mathcal{O}_{i}$ is the localization of $R$ at $\mathfrak{m}_{i}$.
  2. 2.
    $R$ has breadth $n$.
The first point is well-known; see Lemma 3.2.6 and Theorem 3.2.7 in [12]. The second point is [16, Lemma 6.5]. The first point is related to the approximation theorem for valuation topologies [12, Theorem 2.4.1], though note we are not assuming the $\mathcal{O}_{i}$ are independent.
Fact 2.9 ([18, Proposition 2.12]).
Let $R$ be a $W_{n}$-domain. The integral closure $\tilde{R}$ is a multivaluation ring $\bigcap_{i=1}^{m}\mathcal{O}_{i}$ with $m\leq n$.
The width of a poset $(P,\leq)$ is the maximum size of an antichain. Dilworth’s theorem [8] says that if $P$ has width $n<\omega$, then $P$ is a union of $n$ chains. [^4] Dilworth’s theorem is sometimes stated only for finite posets, but it extends to infinite posets by propositional compactness or Tychonoff’s theorem. If $R$ is a ring, let $\operatorname{Spec}R$ denote the poset of prime ideals of $R$ ordered by inclusion.
Remark 2.10.
If $R$ is a ring and $\operatorname{br}(R)\leq n$, then the width of $\operatorname{Spec}R$ is at most $n$. Indeed, suppose $\mathfrak{p}_{1},\ldots,\mathfrak{p}_{n+1}\in\operatorname{Spec}R$ are incomparable. The fact that $\mathfrak{p}_{1}\not\supseteq\mathfrak{p}_{i}$ for $i>1$ implies that $\mathfrak{p}_{1}\not\supseteq\bigcap_{i=2}^{n+1}\mathfrak{p}_{i}$ by [10, Exercise 7.4.11], and so
$\mathfrak{p}_{1}\cap\cdots\cap\mathfrak{p}_{n+1}\neq\mathfrak{p}_{2}\cap\cdots\cap\mathfrak{p}_{n+1}.$
Similarly,
$\mathfrak{p}_{1}\cap\cdots\cap\mathfrak{p}_{n+1}\neq\mathfrak{p}_{1}\cap\cdots\cap\widehat{\mathfrak{p}_{i}}\cap\cdots\cap\mathfrak{p}_{n+1}$
for any $i$. Therefore the lattice of ideals has breadth at least $n+1$, meaning that $\operatorname{br}(R)\geq n+1$.
Fact 2.11.
If $R$ is an NIP ring, then the width of $\operatorname{Spec}R$ is finite [6, Proposition 2.1, Remark 2.2].
3 Henselizing classes
In [22], a certain argument was used to show that NIP $\mathbb{F}_{p}$-algebras satisfy Conjecture 1.2. This same argument can be used in other settings, replacing “NIP $\mathbb{F}_{p}$-algebras” with other classes of NIP rings. In this section we identify the abstract conditions needed to run the argument. Recall that a set $D\subseteq M^{n}$ in a structure $M$ is externally definable if $D=D^{\prime}\cap M^{n}$ for some definable set $D^{\prime}\subseteq N^{n}$ in an elementary extension $N\succeq M$. Equivalently, $D=\phi(M,b)$ for some formula $\phi(x,y)$ and parameter $b$ coming from an elementary extension of $M$. The intuition is that $D$ is defined using parameters from outside $M$. Any prime ideal in a ring is externally definable [7, Fact 4.1].
Remark 3.1.
If $M$ is any structure, the Shelah expansion $M^{\mathrm{Sh}}$ is the expansion of $M$ by all externally definable sets. When $M$ is NIP, the definable sets in $M^{\mathrm{Sh}}$ are exactly the externally definable sets in $M$ [31, Proposition 3.23]. From this, one can show that $M^{\mathrm{Sh}}$ is NIP [31, Corollary 3.24]. A similar argument shows that $M^{\mathrm{Sh}}$ has the same dp-rank as $M$ [7, Fact 3.1]. For example, if $M$ is dp-finite, then $M^{\mathrm{Sh}}$ is dp-finite.
Definition 3.2.
A class $\mathcal{K}$ of rings is pre-henselizing if it is closed under the following conditions:
If $R\in\mathcal{K}$ and $R$ is an integral domain, then any localization $S^{-1}R$ is in $\mathcal{K}$.
If $R\in\mathcal{K}$ and $R^{\prime}\equiv R$, then $R^{\prime}\in\mathcal{K}$.
If $R\in\mathcal{K}$ and $I$ is an externally definable ideal in $R$, then $R/I\in\mathcal{K}$.
Let $R$ be in $\mathcal{K}$, and let $S$ be an $R$-algebra that is free of finite rank as an $R$-module. Then $S\in\mathcal{K}$.
The following classes are pre-henselizing:
  1. 1.
    $\mathbb{F}_{p}$-algebras: the five properties are clear.
  2. 2.
    NIP rings: (El) is well-known. (Loc), (Quot), and (Free) hold because the resulting rings are interpretable in $R^{\mathrm{Sh}}$. For localizations, this is [22, Theorem 2.11].
  3. 3.
    Dp-finite rings: Similar to NIP rings.
Any intersection of pre-henselizing classes is pre-henselizing. For example, the class of NIP $\mathbb{F}_{p}$-algebras is pre-henselizing.
Definition 3.3.
A problematic ring is an integral domain $R$ with exactly two maximal ideals $\mathfrak{m}_{1},\mathfrak{m}_{2}$, such that $R/\mathfrak{m}_{i}$ is infinite for $i=1,2$. A henselizing class of rings is a pre-henselizing class containing no problematic rings.
Lemma 3.9 in [22] says that the class of NIP $\mathbb{F}_{p}$-algebras has no problematic rings, and thus is henselizing.
Proposition 3.4.
Let $\mathcal{K}$ be a henselizing class of NIP rings. Suppose $R\in\mathcal{K}$. Then…
  1. 1.
    $R$ is a finite product of henselian local rings.
If additionally $R$ is an integral domain, then…
  1. 2.
    $R$ is a henselian local ring.
  2. 3.
    The prime ideals of $R$ are linearly ordered by inclusion.
Proof.
The proofs of Lemmas 3.12–3.13, Theorem 3.15, Corollary 3.16, Lemma 3.19, Proposition 3.20, and Theorems 3.21–3.22 in [22] apply with minimal changes, replacing “NIP $\mathbb{F}_{p}$-algebra” with “member of $\mathcal{K}$”. In more detail, if we fix the class $\mathcal{K}$, then the following things are true:
Step 1.
Suppose $R\in\mathcal{K}$ is an integral domain, and $\mathfrak{p}_{1},\mathfrak{p}_{2}$ are two prime ideals such that $R/\mathfrak{p}_{1}$ and $R/\mathfrak{p}_{2}$ are infinite. Then $\mathfrak{p}_{1}$ and $\mathfrak{p}_{2}$ are comparable.
Proof.
Otherwise, let $S=R\setminus(\mathfrak{p}_{1}\cup\mathfrak{p}_{2})$ and let $R^{\prime}$ be $S^{-1}R$, which belongs to $\mathcal{K}$ by (Loc). The ring $R^{\prime}$ has exactly two maximal ideals $\mathfrak{p}_{1}R^{\prime}$ and $\mathfrak{p}_{2}R^{\prime}$, and the quotients $R^{\prime}/\mathfrak{p}_{i}R^{\prime}$ are infinite because of the embeddings $R/\mathfrak{p}_{i}\to R^{\prime}/\mathfrak{p}_{i}R^{\prime}$. Thus $R^{\prime}$ is problematic, contradicting the fact that $\mathcal{K}$ is henselizing. ∎
Step 2.
$\mathcal{K}$ does not contain a domain $R$ with exactly two maximal ideals.
Proof.
The proof of [22, Lemma 3.13] applies, using (El) to make $R$ be saturated at the start of the proof. Step 1 replaces the use of [22, Lemma 3.12]. ∎
Step 3.
If $R\in\mathcal{K}$ is an integral domain, then the prime ideals of $R$ are linearly ordered by inclusion.
Proof.
Otherwise, take two incomparable prime ideals $\mathfrak{p}_{1},\mathfrak{p}_{2}$ in $R$, let $S=\mathfrak{p}_{1}\cup\mathfrak{p}_{2}$, and let $R^{\prime}=S^{-1}R$. Then $R^{\prime}\in\mathcal{K}$ by (Loc), and $R^{\prime}$ has exactly two maximal ideals, contradicting Step 2. ∎
Step 4.
If $R\in\mathcal{K}$ and $\mathfrak{p}_{1},\mathfrak{p}_{2},\mathfrak{q}$ are prime ideals with $\mathfrak{p}_{1},\mathfrak{p}_{2}\supseteq\mathfrak{q}$, then $\mathfrak{p}_{1}$ and $\mathfrak{p}_{2}$ are comparable.
Proof.
Otherwise, $\mathfrak{p}_{1}/\mathfrak{q}$ and $\mathfrak{p}_{2}/\mathfrak{q}$ are incomparable prime ideals in the domain $R/\mathfrak{q}$. But $\mathfrak{q}$ is externally definable in $R$ by [7, Fact 4.1], so $R/\mathfrak{q}\in\mathcal{K}$ by (Quot). This contradicts Step 3. ∎
Recall from the previous section that the width of a poset is the maximum size of an antichain in the poset. As in [22, Definition 3.17], say that a poset $(P,\leq)$ is a forest if $\{x\in P:x\geq c\}$ is a chain for every $c\in P$. Following [22, Definition 3.18], say that a ring $R$ is “good” if the poset of prime ideals in $R$ is a forest of finite width.
Step 5.
If $R\in\mathcal{K}$, then $R$ is good.
Proof.
The poset of prime ideals has finite width by Fact 2.11, because $R$ is NIP. [^5] This is the only point in the proof where we use the assumption that the rings in $\mathcal{K}$ are NIP. It is a forest by Step 4. ∎
Step 6.
If $R\in\mathcal{K}$ is a local ring, then $R$ is henselian.
Proof.
The proof of [22, Proposition 3.20] applies, using (Free) to see that $\mathcal{K}$ contains the ring $R[x_{1},\ldots,x_{n}]/(P_{1}(x_{1}),\ldots,P_{i}(x_{i}))$ appearing in the proof, and using Step 5 to conclude that this ring is “good”. ∎
Step 7.
If $R\in\mathcal{K}$ is arbitrary, then $R$ is a finite product of henselian local rings.
Proof.
$R$ is “good” by Step 5, and therefore a finite product of local rings $R=\prod_{i=1}^{n}R_{i}$ by [22, Lemma 3.19(3)]. The kernel of each projection $R\to R_{i}$ is definable—in fact, generated by an idempotent—and so $R_{i}\in\mathcal{K}$ by (Quot). Then $R_{i}$ is henselian by Step 6. ∎
Step 8.
If $R\in\mathcal{K}$ is integral, then $R$ is a henselian local ring.
Proof.
Clear from Step 7, since $R$ is not a product of more than one ring. ∎
This completes the proof of Proposition 3.4. ∎
As a first application, we mildly strengthen the main theorem of [22]. Recall that the characteristic $\operatorname{char}(R)$ of a ring $R$ is the unique $n\in\mathbb{N}$ such that $n\mathbb{Z}$ is the kernel of the unique homomorphism $\mathbb{Z}\to R$. For example, $\mathbb{Z}/4\mathbb{Z}$ has characteristic 4, though it is not an $\mathbb{F}_{p}$-algebra for any prime $p$.
Theorem 3.5.
Let $R$ be an NIP ring of positive characteristic. Then $R$ is a finite product of henselian local rings.
Proof.
Let $\mathcal{K}_{\mathrm{NIP}}$ be the class of NIP rings, which is pre-henselizing as noted above. Let $\mathcal{K}_{+}$ be the class of rings of positive characteristic. It is easy to see that $\mathcal{K}_{+}$ is pre-henselizing. Then $\mathcal{K}_{\mathrm{NIP}}\cap\mathcal{K}_{+}$, the class of NIP rings of positive characteristic, is pre-henselizing. If $R\in\mathcal{K}_{\mathrm{NIP}}\cap\mathcal{K}_{+}$ is problematic, then $R$ is an integral domain, hence $\operatorname{char}(R)=p$ for some prime $p>0$. Then $R$ is an $\mathbb{F}_{p}$-algebra. But as noted above, there are no problematic NIP $\mathbb{F}_{p}$-algebras, by [22, Lemma 3.9]. Therefore, $\mathcal{K}_{\mathrm{NIP}}\cap\mathcal{K}_{+}$ is henselizing, and Proposition 3.4 applies. ∎
We can use the machinery of Proposition 3.4 to give some reformulations of Conjecture 1.2, the generalized henselianity conjecture:
Proposition 3.6.
The following are equivalent:
  1. 1.
    Conjecture 1.2 holds: every NIP ring is a finite product of henselian local rings.
  2. 2.
    Every NIP integral domain is a henselian local ring.
  3. 3.
    Every NIP integral domain is a local ring.
  4. 4.
    If $R$ is an NIP integral domain, then the prime ideals of $R$ are linearly ordered.
Proof.
$(1)\implies(2)$, $(2)\implies(3)$, and $(4)\implies(3)$ are clear. Suppose (3) holds. Let $\mathcal{K}$ be the class of NIP rings. As noted above, $\mathcal{K}$ is pre-henselizing. By (3), no NIP integral domain has exactly two maximal ideals, which implies that $\mathcal{K}$ is henselizing (Definition 3.3). By Proposition 3.4, every ring in $\mathcal{K}$ is a finite product of henselian local rings, which is (1). By Proposition 3.4 again, the prime ideals in any integral domain in $\mathcal{K}$ are linearly ordered, which is (4). ∎
Remark 3.7.
The generalized henselianity conjecture is meant to be a generalization of Conjecture 1.1, the henselianity conjecture for NIP valuation rings. Arguably, a more direct analogue of Conjecture 1.1 would be the statements:
  1. 5.
    Every NIP local ring is henselian.
  2. 6.
    Every NIP local integral domain is henselian.
In a later paper [24], we will see that (5)–(6) are indeed equivalent to (1)–(4) of Proposition 3.6.
4 NIP rings of finite breadth
If $R$ is a definable ring in some structure, a definable $R$-module means a module $M$ such that the following things are definable in the structure:
  • The underlying set of $M$.
  • The addition map $M\times M\to M$.
  • The scalar multiplication map $R\times M\to M$.
Lemma 4.1.
Let $R$ be a definable ring and $M$ be a definable $R$-module in some structure. Suppose $\operatorname{br}_{R}(M)$ is finite. Then any $R$-submodule $M^{\prime}\subseteq M$ is externally definable.
Proof.
Let $n=\operatorname{br}_{R}(M)<\infty$. The submodule $M^{\prime}$ is a directed union of its finitely generated submodules. Each finitely-generated submodule is generated by at most $n$ elements. In particular, finitely generated submodules of $M$ are uniformly definable. A directed union of uniformly definable sets is externally definable [22, Remark 2.9]. ∎
In this paper, an overring of a domain $R$ is a subring of $\operatorname{Frac}(R)$ containing $R$. For example, the integral closure of $R$ is an overring of $R$.
Corollary 4.2.
If $R$ is a $W_{n}$-domain and $S$ is an overring, then $S$ is externally definable as a subset of $\operatorname{Frac}(R)$, in the structure $R$.
Proof.
If $K=\operatorname{Frac}(R)$, then the breadth of $K$ as an $R$-module equals the breadth of $R$ by [18, Lemma 2.4]. In particular, $K$ has finite breadth, so the $R$-submodule $S$ is externally definable. ∎
For example, the integral closure of $R$ is externally definable. In fact, the integral closure is genuinely definable:
Lemma 4.3.
Let $R$ be a $W_{n}$-domain. Let $\tilde{R}$ be the integral closure of $R$. Then $\tilde{R}=\bigcap_{i=1}^{m}\mathcal{O}_{i}$ for some valuation rings $\mathcal{O}_{i}$. Moreover, $\tilde{R}$ and the $\mathcal{O}_{i}$ are definable subsets of $\operatorname{Frac}(R)$, in the structure $R$.
Proof.
Fact 2.9 gives the decomposition of $\tilde{R}$ as an intersection of valuation rings. For definability of $\tilde{R}$, note that the following are equivalent for $x\in\operatorname{Frac}(R)$:
  1. 1.
    $x$ is integral over $R$.
  2. 2.
    There is an overring $R\subseteq S\subseteq\operatorname{Frac}(R)$ such that $x\in S$ and $S$ is finitely-generated as an $R$-module.
  3. 3.
    There is an overring $R\subseteq S\subseteq\operatorname{Frac}(R)$ such that $x\in S$ and $S$ has the form $Ry_{1}+\cdots Ry_{\ell}$ for some $y_{1},\ldots,y_{\ell}\in\operatorname{Frac}(R)$, where $\ell=\operatorname{br}(R)$.
The equivalence of (1) and (2) is well-known; see [11, Corollary 4.6] for example. The equivalence of (2) and (3) holds because the breadth of $\operatorname{Frac}(R)$ as an $R$-module is at most $n$ by [18, Lemma 2.4]. Condition (3) is a definable property. This shows that $\tilde{R}$ is definable. By Fact 2.8, the $\mathcal{O}_{i}$ are the localizations of $\tilde{R}$ at its maximal ideals. The maximal ideals of $\tilde{R}$ are definable in $\tilde{R}$ by [22, Corollary 2.4], and so each $\mathcal{O}_{i}$ is definable. ∎
Fact 4.4 ([18, Lemma 2.7]).
If $R$ is a $W_{n}$-domain and $S$ is an overring then $S$ is a $W_{n}$-domain.
Remark 4.5.
Let $R$ be a ring and $I$ be an ideal. Then $\operatorname{br}(R/I)\leq\operatorname{br}(R)$, as the lattice of ideals in $R/I$ is a sublattice of the lattice of ideals in $R$.
Remark 4.6.
Let $R$ be a ring and $S$ be an $R$-algebra such that $S\cong R^{d}$ as $R$-modules, for some finite $d$. If $R$ has finite breadth, so does $S$, because
$\operatorname{br}(S)=\operatorname{br}_{S}(S)\leq\operatorname{br}_{R}(S)=\operatorname{br}_{R}(R^{d})=d\cdot\operatorname{br}_{R}(R)=d\operatorname{br}(R)<\infty$
using Facts 2.3 and 2.6.
Lemma 4.7.
Let $\mathcal{K}$ be the class of rings of finite breadth. Then $\mathcal{K}$ is a pre-henselizing class in the sense of Definition 3.2.
Proof.
(Loc) holds by Fact $4.$ (El) is easy, as $\operatorname{br}(R)\leq n$ is expressed by a first-order sentence. (Quot) holds by Remark 4.5, and (Free) holds by Remark 4.6. ∎
For the rest of the section, assume the henselianity conjecture (Conjecture 1.1).
Fact 4.8 (Assuming HC).
Let $\mathcal{O}_{1},\mathcal{O}_{2}$ be two valuation rings on a field $K$. If the structure $(K,\mathcal{O}_{1},\mathcal{O}_{2})$ is NIP, then $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ are comparable.
In fact, Fact 4.8 is equivalent to the henselianity conjecture, by [14, Corollary 5.6].
Lemma 4.9 (Assuming HC).
Let $R$ be an NIP $W_{n}$-domain. Then $\operatorname{Spec}R$ is a chain.
Proof.
Suppose $\mathfrak{p}_{1},\mathfrak{p}_{2}\in\operatorname{Spec}R$ are incomparable. Let $K=\operatorname{Frac}(R)$. By Chevalley’s extension theorem [12, Theorem 3.1.1], there exist valuation rings $(\mathcal{O}_{i},\mathfrak{m}_{i})$ on $K$ for $i=1,2$ such that $\mathcal{O}_{i}\supseteq R$ and $\mathfrak{p}_{i}=\mathfrak{m}_{i}\cap R$. Then $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ are incomparable. (If, say, $\mathcal{O}_{1}\subseteq\mathcal{O}_{2}$, then $\mathfrak{m}_{1}\supseteq\mathfrak{m}_{2}$ and $\mathfrak{p}_{1}\supseteq\mathfrak{p}_{2}$, a contradiction.) By Corollary 4.2, $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ are both definable in $R^{\mathrm{Sh}}$, contradicting Fact 4.8. ∎
In particular, the class of NIP rings of finite breadth is a henselizing class (Definition 3.3), and so Proposition 3.4 applies, giving the following:
Theorem 4.10 (Assuming HC).
Let $R$ be an NIP ring with $\operatorname{br}(R)<\infty$.
  1. 1.
    $R$ is a finite product of henselian local rings.
  2. 2.
    If $R$ is an integral domain, then $R$ is a henselian local domain.
Assuming the henselianity conjecture, we can also strengthen Lemma 4.3:
Lemma 4.11 (Assuming HC).
Let $R$ be an NIP domain with $\operatorname{br}(R)<\infty$. Let $\tilde{R}$ be the integral closure of $R$. Then $\tilde{R}$ is a henselian valuation ring, definable in $R$.
Proof.
By Lemma 4.3, $\tilde{R}$ is definable and an intersection $\bigcap_{i=1}^{n}\mathcal{O}_{i}$ of definable valuation rings, for some $n$. We may assume the $\mathcal{O}_{i}$ are pairwise incomparable. Then $n=1$ by Fact 4.8, and $\mathcal{O}_{1}$ is henselian by the henselianity conjecture. ∎
5 Dp-finite rings
Drop the assumption of the henselianity conjecture.
A definable topology on a structure $M$ means a topology such that some definable family $\{D_{a}\}_{a\in X}$ is a basis of opens. A topology on a field $K$ is a V-topology if it is induced by a valuation or absolute value; see [30, Section 3] and [12, Appendix B] for more about V-topologies. Definable V-topologies play an important role in the conjectural classification of NIP fields [14].
Recall that a structure is dp-finite if its dp-rank is $n$ for some natural number $n$.
Fact 5.1.
  1. 1.
    If $\mathcal{O}$ is a dp-finite valuation ring, then $\mathcal{O}$ is henselian.
  2. 2.
    If $(K,\mathcal{O}_{1},\mathcal{O}_{2})$ is a dp-finite bi-valued field, then $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ are comparable.
Proof.
This is essentially [19, Corollary 4.16], though part (2) requires some explanation. Part (1) is [19, Corollary 4.16(3)]. For part (2), suppose $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ are incomparable. Then $\mathcal{O}_{1}\cdot\mathcal{O}_{2}=\{xy:x\in\mathcal{O}_{1},~y\in\mathcal{O}_{2}\}$ is a definable valuation ring in the structure $(K,\mathcal{O}_{1},\mathcal{O}_{2})$. Let $k$ be the residue field of $\mathcal{O}_{1}\cdot\mathcal{O}_{2}$. The valuation rings $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ induce independent non-trivial valuations on $k$ (see [12], specifically Theorem 2.3.4 and the discussion following Corollary 2.3.2). Replacing $K$ with $k$, we may assume that $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ are independent non-trivial valuations. Then the structure $(K,\mathcal{O}_{1},\mathcal{O}_{2})$ is unstable and has two definable V-topologies, contradicting [19, Corollary 4.16(2)]. ∎
Lemma 5.2.
Let $R$ be a dp-finite $W_{n}$-domain. Then $\operatorname{Spec}R$ is a chain.
Proof.
Suppose $\mathfrak{p}_{1},\mathfrak{p}_{2}\in\operatorname{Spec}R$ are incomparable. As in the proof of Lemma 4.9, [^6] This part of the proof of Lemma 4.9 did not use the assumption of the henselianity conjecture. we can find two incomparable valuation rings $\mathcal{O}_{1},\mathcal{O}_{2}$ on $K=\operatorname{Frac}(R)$ with $\mathcal{O}_{1},\mathcal{O}_{2}\supseteq R$. By Corollary 4.2, $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ are both definable in $R^{\mathrm{Sh}}$, contradicting Fact 5.1(2). ∎
Therefore the class of dp-finite rings of finite breadth is a henselizing class, and Proposition 3.4 yields the following:
Lemma 5.3.
Let $R$ be a dp-finite ring with $\operatorname{br}(R)<\infty$.
  1. 1.
    $R$ is a direct product of finitely many henselian local rings.
  2. 2.
    If $R$ is an integral domain, then $R$ is a henselian local domain and the prime ideals of $R$ are linearly ordered.
We will see shortly that we can remove the finite-breadth assumption in Lemma 5.3.
Lemma 5.4.
Let $N$ be some structure. Let $R$ be a definable ring and $M$ be a definable $R$-module. Let $\operatorname{Sub}^{\mathrm{def}}_{R}(M)$ be the modular lattice of definable $R$-submodules of $M$. Then $\operatorname{br}(M)$ is the breadth of $\operatorname{Sub}^{\mathrm{def}}_{R}(M)$.
Proof.
First of all,
$\operatorname{br}(\operatorname{Sub}^{\mathrm{def}}_{R}(M))\leq\operatorname{br}(\operatorname{Sub}_{R}(M))\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\operatorname{br}(M),$
as $\operatorname{Sub}^{\mathrm{def}}_{R}(M)$ is a sublattice of $\operatorname{Sub}_{R}(M)$. It remains to show that
$n\leq\operatorname{br}(M)\implies n\leq\operatorname{br}(\operatorname{Sub}^{\mathrm{def}}_{R}(M))$
for each finite $n$. Take $x_{1},\ldots,x_{n}\in M$ such that
$Rx_{1}+\cdots+Rx_{n}\neq Rx_{1}+\cdots+\widehat{Rx_{i}}+\cdots+Rx_{n}$
for all $1\leq i\leq n$. Then the definable $R$-submodules $Rx_{1},\ldots,Rx_{n}$ show that $\operatorname{Sub}^{\mathrm{def}}_{R}(M)$ has breadth at least $n$. ∎
Remark 5.5.
If $M$ is a non-trivial $R$-module, then $M$ has a simple subquotient. That is, there are submodules $M^{\prime\prime}<M^{\prime}\leq M$ such that $M^{\prime}/M^{\prime\prime}$ is a simple $R$-module, isomorphic to $R/\mathfrak{m}$ for some maximal ideal $\mathfrak{m}$. To see this, take non-zero $x\in M$, take $M^{\prime}=Rx\subseteq M$, and take $M^{\prime\prime}$ a maximal submodule of $M^{\prime}$ not containing $x$. As a consequence, if $R/\mathfrak{m}$ is infinite for every maximal ideal $\mathfrak{m}$, then every non-trivial $R$-module is infinite.
Lemma 5.6.
Let $R$ be a dp-finite ring. Suppose $R/\mathfrak{m}$ is infinite for every maximal ideal $\mathfrak{m}$ of $R$. Then $\operatorname{br}(R)\leq\operatorname{dp-rk}(R)$. In particular, $\operatorname{br}(R)$ is finite.
Proof.
Suppose $\operatorname{br}(R)\geq n$. By Lemma 5.4, the lattice $\operatorname{Sub}^{\mathrm{def}}_{R}(R)$ of definable ideals has breadth at least $n$, so it contains a sublattice isomorphic to the powerset of $n$. It follows that there are definable ideals $I<I^{\prime}\leq R$ and $J_{1},\ldots,J_{n}$ such that $I<J_{i}\leq I^{\prime}$, and $I^{\prime}/I$ is an internal direct sum $\bigoplus_{i=1}^{n}(J_{i}/I)$. Since dp-rank is additive on products ($\operatorname{dp-rk}(X\times Y)=\operatorname{dp-rk}(X)+\operatorname{dp-rk}(Y)$), the dp-rank of the interpretable group $I^{\prime}/I$ is given by $\operatorname{dp-rk}(I^{\prime}/I)=\sum_{i=1}^{n}\operatorname{dp-rk}(J_{i}/I)$. By Remark 5.5, each interpretable group $J_{i}/I$ is infinite, and so $\operatorname{dp-rk}(J_{i}/I)\geq 1$. Then $\operatorname{dp-rk}(R)\geq\operatorname{dp-rk}(I^{\prime})\geq\operatorname{dp-rk}(I^{\prime}/I)=\sum_{i=1}^{n}\operatorname{dp-rk}(J_{i}/I)\geq n$. We have shown
$n\leq\operatorname{br}(R)\implies n\leq\operatorname{dp-rk}(R).\qed$
Example 5.7.
If $(R,\mathfrak{m})$ is a dp-minimal local domain and the residue field $R/\mathfrak{m}$ is infinite, then $\operatorname{br}(R)\leq\operatorname{dp-rk}(R)=1$, so $R$ is a $W_{1}$-domain, or equivalently, a valuation ring. This was originally proven by d’Elbée and Halevi [6, Theorem 4.1].
Lemma 5.8.
Let $R$ be a problematic ring (Definition 3.3), i.e., a domain with exactly two maximal ideals $\mathfrak{m}_{1},\mathfrak{m}_{2}$, such that the quotients $R/\mathfrak{m}_{1}$ and $R/\mathfrak{m}_{2}$ are infinite. Then $R$ is not dp-finite.
Proof.
Suppose otherwise. By Lemma 5.6, $\operatorname{br}(R)$ is finite. The incomparability of $\mathfrak{m}_{1}$ and $\mathfrak{m}_{2}$ contradicts Lemma 5.2. ∎
It follows that the class of dp-finite rings is henselizing. Then Proposition 3.4 gives the following:
Theorem 5.9.
Let $R$ be a dp-finite ring.
  1. 1.
    $R$ is a direct product of finitely many henselian local rings.
  2. 2.
    If $R$ is an integral domain, then $R$ is a henselian local domain and the prime ideals of $R$ are linearly ordered.
In the dp-minimal case, the fact that the prime ideals are linearly ordered is due to d’Elbée and Halevi [6, Corollary 2.4].
6 Dp-finite $W_{n}$-domains
Before focusing on Noetherian rings, there are a few more results we can say about dp-finite rings of finite breadth.
Proposition 6.1.
Let $R$ be a dp-finite $W_{n}$-domain with fraction field $K$. Suppose $R\neq K$. Let $\tilde{R}$ be the integral closure of $R$.
  1. 1.
    $\tilde{R}$ is a henselian valuation ring, definable in the structure $R$ as a subset of $\operatorname{Frac}(R)$.
  2. 2.
[…]
Recall that $R$ is local with maximal ideal $\mathfrak{m}$. Let $\tilde{\mathfrak{m}}$ denote the maximal ideal of $\tilde{R}$.
  1. 3.
    The inclusion $R\to\tilde{R}$ is a local homomorphism, meaning that $\mathfrak{m}=R\cap\tilde{\mathfrak{m}}$.
  2. 4.
    $R$ and $\tilde{R}$ have the same residue characteristic.
  3. 5.
    $R/\mathfrak{m}$ is finite iff $\tilde{R}/\tilde{\mathfrak{m}}$ is finite. More generally, $\tilde{R}/\tilde{\mathfrak{m}}$ is a finite extension of $R/\mathfrak{m}$.
Proof.
  1. 1.
    Lemma 4.3 shows that $\tilde{R}$ is a finite intersection of definable valuation rings $\bigcap_{i=1}^{m}\mathcal{O}_{i}$. By Fact 5.1(1), the $\mathcal{O}_{i}$ are henselian. By Fact 5.1(2), the $\mathcal{O}_{i}$ are pairwise comparable, so we can take $m=1$.
  2. 2.
[…]
  3. 3.
    Suppose $x\in R$. We must show $x\in\mathfrak{m}\Leftrightarrow x\in\tilde{\mathfrak{m}}$. One direction is easy: if $x\notin\mathfrak{m}$ then $x\in R^{\times}\subseteq\tilde{R}^{\times}$, so $x\notin\tilde{\mathfrak{m}}$. Conversely, suppose $x\notin\tilde{\mathfrak{m}}$ but $x\in\mathfrak{m}$. As $\tilde{R}$ is a valuation ring, $1/x\in\tilde{R}$, meaning that $1/x$ is integral over $R$. Therefore there are $c_{0},\ldots,c_{n-1}\in R$ such that $x^{-n}=c_{0}+c_{1}x^{-1}+\cdots+c_{n-1}x^{1-n}$. Then $1=c_{0}x^{n}+\cdots+c_{n-1}x\in\mathfrak{m}$, which is absurd.
  4. 4.
    Part (3) gives a field embedding $R/\mathfrak{m}\to\tilde{R}/\tilde{\mathfrak{m}}$, so these must have the same characteristic.
  5. 5.
    The $(\Leftarrow)$ direction follows from the field embedding $R/\mathfrak{m}\to\tilde{R}/\tilde{\mathfrak{m}}$. Conversely, suppose $R/\mathfrak{m}$ is finite. Note that $\tilde{R}$ and $\tilde{\mathfrak{m}}$ are $\tilde{R}-$submodules of $K$, hence $R$-submodules of $K$ as $R\subseteq\tilde{R}$. By [18, Lemma 2.4], $\operatorname{br}_{R}(K)=\operatorname{br}_{R}(R)<\infty$, and so
    $\operatorname{br}_{R}(\tilde{R}/\tilde{\mathfrak{m}})\leq\operatorname{br}_{R}(\tilde{R})\leq\operatorname{br}_{R}(K)=\operatorname{br}(R)<\infty$
    by Fact 2.6. We claim that $\mathfrak{m}$ annihilates the $R$-module $\tilde{R}/\tilde{\mathfrak{m}}$. Indeed, if $x\in\mathfrak{m}$ and $y\in\tilde{R}$, then $x\in\tilde{\mathfrak{m}}$ by part (3), and so $xy\in\tilde{\mathfrak{m}}\tilde{R}=\tilde{\mathfrak{m}}$. Therefore $\tilde{R}/\tilde{\mathfrak{m}}$ is an $R/\mathfrak{m}$-module, and
[…]
    by Remarks 2.4 and 2.5. Therefore if $R/\mathfrak{m}$ is finite then so is $\tilde{R}/\tilde{\mathfrak{m}}$. ∎
Corollary 6.2.
Let $R$ be a dp-finite $W_{n}$-domain with $R\neq\operatorname{Frac}(R)$. If $\operatorname{Frac}(R)$ has positive characteristic, then the residue field of $R$ is infinite and $\{x^{p}-x:x\in R\}=R$.
Proof.
Let $\tilde{R}$ be the integral closure. Then $\tilde{R}$ is a positive characteristic NIP valuation ring, so it has infinite residue field by [26, Proposition 5.3]. By Proposition 6.1, $R$ has infinite residue field, too. The surjectivity of the Artin-Schreier map $R\to R$ then follows by [22, Theorem 3.4]. ∎
7 NIP Noetherian rings
Lemma 7.1.
Let $R$ be a ring. Then $\operatorname{br}(R)\geq n$ iff there are ideals $I\leq I^{\prime}\leq R$ such that the $R$-module $I^{\prime}/I$ is a direct sum of $n$ simple $R$-modules.
Proof.
The condition certainly implies $\operatorname{br}(R)\geq n$. Conversely, supose $\operatorname{br}(R)\geq n$. Then there are ideals $I_{0}\leq I^{\prime}_{0}\leq R$ such that the $R$-module $I^{\prime}_{0}/I_{0}$ is a direct sum of $n$ non-trivial $R$-modules.
$I^{\prime}_{0}/I_{0}\cong\bigoplus_{i=1}^{n}N_{i}.$
Every non-trivial $R$-module has a simple subquotient (see Remark 5.5), so there are submodules $N^{\prime}_{i}\leq N^{\prime\prime}_{i}\leq N_{i}$ such that $N^{\prime\prime}_{i}/N^{\prime}_{i}$ is simple. Then
$\left.\bigoplus_{i=1}^{n}N^{\prime\prime}_{i}\middle/\bigoplus_{i=1}^{n}N^{\prime}_{i}\right.\cong\bigoplus_{i=1}^{n}N^{\prime\prime}_{i}/N^{\prime}_{i}$
is a subquotient of $I^{\prime}_{0}/I_{0}$ isomorphic to a direct sum of $n$ simple $R$-modules. A subquotient of $I^{\prime}_{0}/I_{0}$ has the form $I^{\prime}/I$ for some ideals $I,I^{\prime}$ with $I_{0}\leq I\leq I^{\prime}\leq I^{\prime}_{0}$. ∎
If $R$ is a Noetherian ring, let $\dim(R)$ denote the Krull dimension of $R$. Recall that if $R$ is a Noetherian ring of Krull dimension 0, then $R$ has finite length (as an $R$-module) [11, Theorem 2.14]. A ring $R$ is semilocal if it has finitely many maximal ideals [11, Exercise 4.13].
Lemma 7.2.
Let $R$ be a semilocal Noetherian domain with $\dim(R)\leq 1$. Then $\operatorname{br}(R)$ is finite.
Proof.
If $\dim(R)=0$, then $R$ is a field, and $\operatorname{br}(R)=1$. Suppose $\dim(R)=1$. Let $\mathfrak{m}_{1},\ldots,\mathfrak{m}_{n}$ be the maximal ideals of $R$. The only other prime ideal is the minimal prime ideal 0.
Note that if $I$ is any ideal in $R$, then the poset $\operatorname{Spec}R/I$ is isomorphic to the subposet $\{\mathfrak{p}\in\operatorname{Spec}R:\mathfrak{p}\supseteq I\}$ of $\operatorname{Spec}R$. In the case when $I\neq 0$, this poset is a subset of the antichain $\{\mathfrak{m}_{1},\ldots,\mathfrak{m}_{n}\}$, so $\dim(R/I)=0$ and $R/I$ has finite length. Letting $\ell_{R}(M)$ denote the length of an $R$-module $M$, we have
$I\neq 0\implies\ell_{R}(R/I)=\ell_{R/I}(R/I)<\infty.$ ($\ast$)
Let $J=\bigcap_{i=1}^{n}\mathfrak{m}_{i}$ (the Jacobson radical); it is non-zero because an intersection of two non-zero ideals is non-zero in a domain. Take non-zero $x\in J$. By ($\ast$), the $R$-module $R/xR$ has finite length $m$. We claim $\operatorname{br}(R)\leq m$. Otherwise by Remark 7.1 there are ideals $I^{\prime}<I\leq R$ such that $I/I^{\prime}$ is a direct sum of $m+1$ simple $R$-modules. Recall that any simple $R$-module $M$ is isomorphic to $R/\mathfrak{m}$ for some maximal ideal $\mathfrak{m}$, and so the annihilator $\operatorname{Ann}(M)=\mathfrak{m}$ is a maximal ideal. Because $x$ is in every maximal ideal, it annihilates any simple $R$-module, and therefore it annihilates $I/I^{\prime}$, meaning that $xI\leq I^{\prime}$. Then
$m+1=\ell_{R}(I/I^{\prime})\leq\ell_{R}(I/xI).$
By ($\ast$), the modules $R/I$ and $R/xI$ have finite length, and so
$\ell_{R}(R/I)+\ell_{R}(I/xI)=\ell_{R}(R/xI)=\ell_{R}(R/xR)+\ell_{R}(xR/xI)<\infty$
by the additivity of length (Fact 2.7). Multiplication by $x$ gives an isomorphism of $R$-modules from $R/I$ to $xR/xI$, so $\ell_{R}(R/I)=\ell_{R}(xR/xI)$. Canceling this from both sides, we see $\ell_{R}(I/xI)=\ell_{R}(R/xR)=m$. Then
$m+1=\ell_{R}(I/I^{\prime})\leq\ell_{R}(I/xI)=m,$
a contradiction. ∎
Lemma 7.3.
Let $R$ be a semilocal Noetherian ring with $\dim(R)\leq 1$. Then $\operatorname{br}(R)$ is finite.
Proof.
Recall that Noetherian rings have only finitely many minimal prime ideals [11, Section 3.1]. Let $\mathfrak{p}_{1},\ldots,\mathfrak{p}_{n}$ denote the finitely many minimal prime ideals of $R$. By Lemma 7.2, the ring $R/\mathfrak{p}_{i}$ has finite breadth. By Fact 2.6, it follows that any finitely-generated $R/\mathfrak{p}_{i}$-module has finite breadth as an $R/\mathfrak{p}_{i}$-module, or equivalently, as an $R$-module.
By a basic fact in commutative algebra [11, Corollary 2.12], the set of nilpotent elements (i.e., the nilradical $\sqrt{0}$) is exactly $\bigcap_{i=1}^{n}\mathfrak{p}_{i}$. In particular, the product $I=\mathfrak{p}_{1}\mathfrak{p}_{2}\cdots\mathfrak{p}_{n}$ is contained in the nilradical—every element of $I$ is nilpotent. As $I$ is finitely-generated, it follows that $I^{k}=0$ for some $k$. Thus $\mathfrak{p}_{1}^{k}\mathfrak{p}_{2}^{k}\cdots\mathfrak{p}_{n}^{k}=0$ for some $k$. Therefore there is a finite sequence $\mathfrak{q}_{1},\ldots,\mathfrak{q}_{m}$ such that each $\mathfrak{q}_{i}$ is a minimal prime, and $\mathfrak{q}_{1}\mathfrak{q}_{2}\cdots\mathfrak{q}_{m}=0$. Let $I_{j}=\mathfrak{q}_{1}\cdots\mathfrak{q}_{j}$ for $0\leq j\leq m$. This gives a descending sequence of ideals
$R=I_{0}\geq I_{1}\geq\cdots\geq I_{N}=0.$
For each $j>0$, $I_{j-1}/I_{j}=I_{j-1}/\mathfrak{q}_{j}I_{j-1}$ is an $R/\mathfrak{q}_{j}$-module, finitely generated by Noetherianity. By the first paragraph of the proof, $I_{j-1}/I_{j}$ has finite breadth as an $R$-module. By Fact 2.6 again, $R=I_{0}/I_{N}$ has finite breadth, bounded by $\sum_{j=1}^{N}\operatorname{br}(I_{j}/I_{j-1})$. ∎
Remark 7.4.
Lemma 7.3 is related to—and possibly a consequence of—work of I. S. Cohen. In [5, Section 4], Cohen says that a ring $R$ has “rank kk” if every ideal $I\subseteq R$ is generated by $k$ or fewer generators. This condition is at first glance similar to the condition $\operatorname{br}(R)\leq k$, which means that any finite set of generators for an ideal can be shrunk to a set of size $k$. In general the two conditions are orthogonal. For example, $\mathbb{Z}$ has infinite breadth, but “rank 1” in the sense of Cohen. Conversely, if $R$ is a non-discrete valuation ring, then $R$ has breadth 1, but does not have “finite rank” in the sense of Cohen, as $R$ is non-Noetherian. When $R$ is a Noetherian ring, $\operatorname{br}(R)\leq k$ implies Cohen’s “rank kk”. In fact, when $(R,\mathfrak{m})$ is a local Noetherian ring, “rank kk” is equivalent to $\operatorname{br}(R)\leq k$” because Nakayama’s lemma makes both conditions equivalent to the condition
$\dim_{R/\mathfrak{m}}(I/\mathfrak{m}I)\leq k\text{ for any ideal }I\subseteq R.$
Cohen shows that if $R$ is a local Noetherian domain, then $R$ has finite rank if and only if $\dim(R)\leq 1$ [5, Theorem 9] (see also [5, Corollary 1 to Theorem 1]). This gives another proof of Lemma 7.3, at least in the case when $R$ is a local domain. (The direction of Cohen’s result that is relevant to Lemma 7.3 may go back to [1].)
Proposition 7.5.
Let $R$ be a Noetherian ring. The following are equivalent:
  1. 1.
    $\operatorname{br}(R)<\infty$.
  2. 2.
    The poset $\operatorname{Spec}R$ has finite width.
  3. 3.
    $\operatorname{Spec}R$ is finite.
  4. 4.
    $R$ is semilocal with $\dim(R)\leq 1$.
Proof.
$(1)\implies(2).$
Remark 2.10.
$(2)\implies(3).$
By Dilworth’s theorem [8], $\operatorname{Spec}R$ is a finite union of chains. In a Noetherian ring, the poset $\operatorname{Spec}R$ satisfies the ascending chain condition (by Noetherianity) and the descending chain condition (by dimension theory; see [11, Corollary 10.3]). Therefore each chain in the union is finite.
$(3)\implies(4).$
The fact that $R$ is semilocal is clear. Suppose for the sake of contradiction that $\dim(R)\geq 2$. Take three prime ideals $\mathfrak{p}_{1}\subsetneq\mathfrak{p}_{2}\subsetneq\mathfrak{p}_{3}$. By [27, Theorem 144], the fact that there is a prime ideal between $\mathfrak{p}_{1}$ and $\mathfrak{p}_{3}$ implies that there are infinitely many such prime ideals. This contradicts (3).
$(4)\implies(1).$
Lemma 7.3. ∎
Corollary 7.6.
Let $R$ be an NIP Noetherian ring. Then $R$ has finite breadth, $R$ is semilocal with $\dim(R)\leq 1$, and $\operatorname{Spec}R$ is finite.
Proof.
If $R$ is an NIP ring, then $\operatorname{Spec}R$ has finite width by Fact 2.11. ∎
Lemma 7.7.
Let $R$ be an NIP Noetherian domain with fraction field $K$. Let $\tilde{R}$ be the integral closure of $R$ (in $K$).
  1. 1.
    $\tilde{R}$ is a finite intersection of discrete valuation rings, each of which is definable in $R$.
  2. 2.
    If $R$ is dp-finite or the henselianity conjecture holds, then $\tilde{R}$ is a henselian DVR.
Proof.
By Corollary 7.6, $R$ has finite breadth and $\dim(R)\leq 1$.
  1. 1.
    By Lemma 4.3, $\tilde{R}$ is an intersection of definable valuation rings $\mathcal{O}_{1}\cap\cdots\cap\mathcal{O}_{n}$. Because $\dim(R)\leq 1$, the Krull-Akizuki theorem [11, Theorem 11.13] shows that all overrings of $R$ are Noetherian and 1-dimensional. Therefore each valuation ring $\mathcal{O}_{i}$ is Noetherian, and therefore a discrete valuation ring.
  2. 2.
    By Proposition 6.1(1) or Lemma 4.11, $\tilde{R}$ is a henselian valuation ring. ∎
Theorem 7.8.
Let $R$ be an NIP Noetherian domain. If $R$ is not a field, then $\operatorname{Frac}(R)$ has characteristic 0.
Proof.
By Lemma 7.7, the integral closure $\tilde{R}$ is an intersection of discrete valuation rings $\bigcap_{i=1}^{m}\mathcal{O}_{i}$, with each $\mathcal{O}_{i}$ definable in the structure $R$. If $R\neq K$, then $\tilde{R}\neq K$. (Take a non-zero maximal ideal $\mathfrak{m}\in\operatorname{Spec}R$. Take $x\in\mathfrak{m}$. Then $1/x$ is not integral over $R$, or else $x^{-n}+c_{n-1}x^{1-n}+\cdots+c_{1}x^{-1}+c_{0}=0$ for some $c_{i}$ in $R$. This would imply $1=c_{n-1}x+c_{n-2}x^{2}+\cdots+c_{0}x^{n}\in\mathfrak{m}$, which is absurd.) Therefore some $\mathcal{O}_{i}$ is not KK. [^7] More precisely, since $K$ is not a DVR, the claim is that $m>0$. Then there is at least one DVR $\mathcal{O}$ on $K$, definable in the structure $R$. By [26, Proposition 5.4], there are no NIP discrete valuation rings of positive characteristic, so $\operatorname{char}(K)=0$. ∎
Now assume the henselianity conjecture (Conjecture 1.1).
Theorem 7.9 (Assuming HC).
Let $R$ be an NIP Noetherian ring.
  1. 1.
    $R$ is a finite product of henselian local rings.
  2. 2.
    If $R$ is an integral domain, then $R$ is a henselian local domain.
Proof.
This follows directly from Theorem 4.10, since we know $\operatorname{br}(R)<\infty$ by Corollary 7.6. ∎
7.1 Rings elementarily equivalent to NIP Noetherian rings
Drop the assumption of the henselianity conjecture.
Proposition 7.10.
Let $R$ be an NIP Noetherian ring. The family of ideals in $R$ is uniformly definable.
Proof.
Let $n=\operatorname{br}(R)<\infty$. By Noetherianity, every ideal is finitely generated. Then every ideal is generated by $n$ or fewer generators. ∎
Proposition 7.11.
Let $R_{0}$ be an NIP Noetherian ring, and suppose $R\equiv R_{0}$.
  1. 1.
    $R$ has finite breadth $n<\infty$.
  2. 2.
    An ideal $I\subseteq R$ is definable if and only if it is finitely generated if and only if it is generated by $n$ or fewer generators.
  3. 3.
    Any ideal in $R$ is externally definable.
  4. 4.
    The family $\mathcal{I}$ of definable ideals in $R$ is uniformly definable.
  5. 5.
    If $\mathcal{F}$ is a non-empty definable family of ideals, then $\mathcal{F}$ has a maximal element.
Proof.
The condition $\operatorname{br}(R)\leq n$” is expressed by a first-order sentence, so breadth is preserved under elementary equivalence. Therefore (1) holds by Corollary 7.6. Then (3) holds by Lemma 4.1.
For (2), the fact that $\operatorname{br}(R)=n$ shows that every finitely generated ideal is generated by $n$ elements. Finitely generated ideals are certainly definable. Conversely, suppose $I\subseteq R$ is definable, defined by a formula $\phi(x,b)$ with parameters $b\in I$. As in the proof of Proposition 7.10, the ring $R_{0}$ satisfies the following property:
For any $b^{\prime}$, if $\phi(R_{0},b^{\prime})$ is an ideal, then $\phi(R_{0},b^{\prime})$ is generated by $n$ elements.
This is expressed by a first-order sentence, so it holds in $R$, and therefore $\phi(R,b)$ is generated by $n$ elements.
Once (2) is known, (4) follows immediately. For part (5), note that the ring $R_{0}$ satisfies the following property, by Noetherianity:
Let $\mathcal{I}_{n}$ be the family of ideals generated by $n$ elements. If $\mathcal{F}$ is a non-empty definable subfamily of $\mathcal{I}_{n}$, then $\mathcal{F}$ has a maximal element.
This can be expressed by an axiom schema, so it holds in $R$. But in $R$, every definable ideal is generated by $n$ elements, by part (2). Therefore (5) holds. ∎
Proposition 7.12.
Let $R$ be a Noetherian ring and $R_{0}$ be an elementary substructure. Then $R_{0}$ is Noetherian.
Proof.
Otherwise, take an ideal $I\subseteq R_{0}$ that is not finitely generated. Recursively choose $x_{1},x_{2},\ldots\in I$ by taking $x_{n}\in I\setminus(x_{1},\ldots,x_{n-1})\neq\varnothing$. Then we have an ascending chain of finitely-generated ideals in $R_{0}$:
$0\subsetneq(x_{1})\subsetneq(x_{1},x_{2})\subsetneq\cdots$
In the ring $R$, the chain
$0\subseteq(x_{1})_{R}\subseteq(x_{1},x_{2})_{R}\subseteq\cdots$
cannot be strictly ascending, because $R$ is Noetherian. Therefore there is some $n$ such that $x_{n}\in(x_{1},\ldots,x_{n-1})_{R}=Rx_{1}+\cdots+Rx_{n-1}$. As $R_{0}\preceq R$, we also have $x_{n}\in R_{0}x_{1}+\cdots+R_{0}x_{n-1}$, contradicting the choice of $x_{n}$. ∎
8 Dp-finite Noetherian domains
Fact 8.1 (part of [23, Theorem 2.11]).
Let $(\mathcal{O},\mathfrak{m})$ be a dp-finite (or strongly dependent) non-trivially valued field with residue field $k$ and value group $\Gamma$. Suppose $\operatorname{char}(k)=p>0$. Then one of three things happens:
  1. 1.
    $\operatorname{char}(K)=0$, $k$ is finite, and the interval $v(p),v(p)]\subseteq\Gamma$ is finite.
  2. 2.
    $\operatorname{char}(K)=0$, $k$ is infinite, and the interval $v(p),v(p)]\subseteq\Gamma$ is $p$-divisible.
  3. 3.
    $\operatorname{char}(K)=p$, $k$ is infinite, and $\Gamma$ is $p$-divisible.
Let $R$ be a dp-finite Noetherian domain with $R\neq\operatorname{Frac}(R)=K$. Let $\tilde{R}$ be the integral closure of $R$. Here is what we know so far:
  • $\operatorname{char}(K)=0$ (Theorem 7.8).
  • $R$ is a henselian local domain (Theorem 5.9).
  • $R$ has Krull dimension 1 and finite breadth (Corollary 7.6).
  • $\tilde{R}$ is a definable henselian DVR (Lemma 7.7).
[…]
  • $R$ and $\tilde{R}$ have the same residue characteristic (Proposition 6.1(4)).
  • The residue field of $R$ is finite iff the residue field of $\tilde{R}$ is finite (Proposition 6.1(5)).
As $\tilde{R}$ is definable, it is also dp-finite. By Fact 8.1 and the fact that $\tilde{R}$ is a discrete valuation ring, one of two things happens:
  1. 1.
    $\tilde{R}$ has residue characteristic 0.
  2. 2.
    $\tilde{R}$ has finite residue field.
(Cases (2)–(3) of Fact 8.1 cannot happen because the valuation is discrete.) By parts (4) and (5) of Proposition 6.1, the properties of the residue field of $\tilde{R}$ transfer back to the residue field of $R$. Putting everything together, we get the following trichotomy for dp-finite Noetherian domains:
Theorem 8.2.
Let $R$ be a dp-finite Noetherian domain with fraction field $K$. Recall that $R$ is local; let $k$ be the residue field. Then one of three things happpens:
  1. 1.
    $R$ is a field.
  2. 2.
    $R$ is not a field. $K$ and $k$ have characteristic 0.
  3. 3.
    $R$ is not a field. $K$ has characteristic 0 and $k$ is finite.
Remark 8.3.
Theorem 8.2 does not hold for general NIP Noetherian rings. Let $\mathbb{Q}_{p}^{un}$ be the maximal unramified algebraic extension of $\mathbb{Q}_{p}$. Then $\mathbb{Q}_{p}^{un}$ is NIP [3, Corollaire 7.5]. If $R$ is the ring of integers of $\mathbb{Q}_{p}^{un}$, then $R$ does not satisfy Theorem 8.2, because the residue field $\mathbb{F}_{p}^{alg}$ is infinite with positive characteristic. The field $\mathbb{Q}_{p}^{un}$ and ring $R$ are not dp-finite, or even strongly dependent by Fact 8.1.
8.1 The classification of dp-minimal Noetherian domains
If $R$ is a ring and $A,B\subseteq R$, let
$A+B=\{x+y:x\in A,~y\in B\}$
$A-B=\{x-y:x\in A,~y\in B\}.$
In particular, ABA-B does not mean $A\setminus B$. Recall the definition of “definable V-topology” from the start of Section 5.
Fact 8.4 ([15, Theorem 1.3]).
Let $(K,+,\cdot,\ldots)$ be an expansion of a field. Suppose $K$ is dp-minimal but not strongly minimal. Then there is a non-discrete, Hausdorff, definable field topology $\tau_{K}$ on $K$ characterized by the fact that the following family is a neighborhood basis of 0:
$\{X-X:X\subseteq K\text{ is definable and infinite}\}.$
Moreover, $\tau_{K}$ is a V-topology.
The topology $\tau_{K}$ is called the canonical topology on $(K,+,\cdot,\ldots)$.
Lemma 8.5.
Let $(K,+,\cdot,\ldots)$ be as in Fact 8.4. Let $\tau_{0}$ be a non-discrete, Hausdorff, definable ring topology on $K$. Then $\tau_{0}$ is the canonical topology $\tau_{K}$.
Proof.
Among non-discrete Hausdorff ring topologies on $K$, V-topologies are maximal with respect to coarsening—there are no strictly coarser ring topologies [30, Theorem 3.2]. As $\tau_{K}$ is a V-topology, it suffices to show that $\tau_{0}$ is coarser than $\tau_{K}$. Let $U$ be a $\tau_{0}$-neighborhood of 0. It suffices to show that $U$ is a $\tau_{K}$-neighborhood of 0. By $\tau_{0}$-continuity of +,+,-, there is a $\tau_{0}$-neighborhood $V$ of 0 such that $V-V\subseteq U$. As $\tau_{0}$ is definable, we can take $V$ to be a definable set. As $\tau_{0}$ is a non-discrete ring topology, $0\in K$ is not $\tau_{0}$-isolated. Therefore $V$ is infinite. Then VVV-V is a $\tau_{K}$-neighborhood of 0 by definition of $\tau_{K}$, and so $U$ is a $\tau_{K}$-neighborhood of 0. ∎
If $R\neq K=\operatorname{Frac}(R)$, then $R$ induces a ring topology on $K$ characterized by the fact that $\{aR:a\in K^{\times}\}$ is a neighborhood basis of 0, or equivalently, characterized by the fact that the set of non-zero ideals of $R$ form a neighborhood basis of 0 [30, Example 1.2]. Following [25], we call this topology the $R$-adic topology.
Lemma 8.6.
Suppose $R\neq K=\operatorname{Frac}(R)$ and $R$ is dp-minimal and Noetherian. Let $\tilde{R}$ be the integral closure of $R$. Then the $R$-adic topology agrees with the $\tilde{R}-$adic topology. Moreover there is $a\in K^{\times}$ such that $R\supseteq a\tilde{R}$.
Proof.
[…]
Recall the classification of dp-minimal DVRs mentioned in the introduction (Fact 1.12), including the division into equicharacteristic and mixed characteristic cases. Note that the equicharacteristic cases always have characteristic 0.
Lemma 8.7.
Let $R$ be a mixed characteristic dp-minimal DVR. If $a\in R\setminus\{0\}$, then $R/aR$ is finite.
Proof.
Let $\pi\in R$ generate the maximal ideal. Then $a=\pi^{n}u$ where $n$ is the valuation of $a$ and $u$ is a unit, and so $R/aR=R/\pi^{n}R$. We claim that $R/\pi^{n}R$ is finite by induction on $n$. The base case $n=1$ holds by Fact 1.12. If $n>0$, then we have a short exact sequence
$0\to\pi R/\pi^{n}R\to R/\pi^{n}R\to R/\pi R\to 0$
and an isomorphism $\pi R/\pi^{n}R\to R/\pi^{n-1}R$. By induction, $R/\pi R$ and $R/\pi^{n-1}R$ are finite, so $R/\pi^{n}R$ is finite. ∎
If $R$ is a ring, a finite-index subring is a subring $S\subseteq R$ with finite index in the additive group $(R,+)$.
Lemma 8.8.
Suppose $R$ is a mixed characteristic dp-minimal DVR and $S$ is a finite index subring.
  1. 1.
    $S$ is dp-minimal.
  2. 2.
    $S$ is Noetherian.
Proof.
  1. 1.
    It suffices to show that $S$ is definable in $R$. Let $n$ be the index of $S$ in $R$. Then $n$ annihilates $R/S$, so $S\supseteq nR$. As $\operatorname{Frac}(R)$ has characteristic 0, $R\models n\neq 0$ and so $R/nR$ is finite by Lemma 8.7. Then $S$ is a finite union of additive cosets of $nR$, so $S$ is definable.
  2. 2.
    As in the previous point, $S\supseteq nR$ for some $n$. Let $I$ be an ideal in $S$. We claim that $I$ is finitely generated. We may assume $I\neq 0$. Take non-zero $a\in I\subseteq S\subseteq R$. Then
    $anR\subseteq aS\subseteq I\subseteq S\subseteq R.$
    By Lemma 8.7, $R/anR$ is finite. Therefore $I/aS$ is finite. Let $\{b_{1},\ldots,b_{m}\}$ be a finite set of coset representatives of $aS$ in $I$. Then $I$ is generated as an $S$-ideal by $\{a,b_{1},\ldots,b_{m}\}$. ∎
We can now classify dp-minimal Noetherian domains.
Theorem 8.9.
The dp-minimal Noetherian domains are precisely the following:
  1. 1.
    Dp-minimal fields.
  2. 2.
    Equicharacteristic dp-minimal DVRs.
  3. 3.
    Finite-index subrings of mixed characteristic dp-minimal DVRs.
Proof.
The three cases are all dp-minimal Noetherian domains. For (1) and (2) this is obvious, and for (3) this is Lemma 8.8. Conversely, suppose $R$ is a dp-minimal integral domain. We claim that $R$ falls into one of the three cases. By Theorem 5.9, $R$ is a local ring. By Theorem 8.2, one of three things happens:
  • $R$ is a field. This is case (1).
  • $R$ is not a field; the fraction field and residue field both have characteristic 0. Then the residue field is infinite, and so $R$ is a valuation ring by Example 5.7.
  • $R$ is not a field; the fraction field has characteristic 0 and the residue field is finite. By Proposition 6.1(2) and Lemma 7.7, the integral closure $\tilde{R}$ is a dp-minimal DVR. By Proposition 6.1(4), $\tilde{R}$ is mixed characteristic. By Lemma 8.6, there is non-zero $a$ such that $R\supseteq a\tilde{R}$. Note
    $a\in a\tilde{R}\subseteq R\subseteq\tilde{R}$,
    so $a$ is a non-zero element of $\tilde{R}$. Then $\tilde{R}/a\tilde{R}$ is finite by Lemma 8.7, implying that $\tilde{R}/R$ is finite. ∎
Acknowledgments.
The author was supported by the National Natural Science Foundation of China (Grant No. 12101131). The author would like to thank two anonymous referees who found a number of errors and offered many helpful comments.